### v2

@@ Abstract
This paper integrates a novel concept, coined as an individual–Environmental, Social, Economics (ESE) score with a joint liability model in order to gain insights into optimal group sizes and individual–ESE score relationships. We adopt a mean–variance utility function for farmers to effectively capture risk associated with anticipated profits. Through a set of simulation exercises, the paper studies a number of implications of incorporating $ESE$ scores into credit evaluation systems, offering, in the process, a nuanced comprehension of the repercussions under various climate conditions.
@@ 1 Introduction
In recent years the importance of incorporating sustainability factors into financial decision–making has gained a major traction. At the same time, while credit metrics for corporations have been extensively studied, there remains a noticeable gap in the assessment of individual borrowers, particularly farmers seeking credit to finance their agricultural operations. To address this gap, this paper introduces a novel concept of scoring system coined as an Environmental, Sustainability, and Economics (ESE), which considers not only traditional credit metrics, including economic factors, but importantly also environmental and social aspects which contribute to sustainable agricultural practices and which are so prominent in the standard Environmental, Sustainability, and Governance (ESG) score. See ztwp2024 for a further discussion and analyis of this individual–ESE concept.
@@
In particular, in this paper, we propose a novel link between an individual–ESE score and a joint liability model and provide lenders and borrowers (where lenders are insurance companies and borrowers are farmers) with optimal scores and corresponding optimal group sizes by maximizing total utility of a group of individual borrowers. We start our analysis with the simplest case of two borrowers in the group. Later we generalize it to a group of $n$ individual borrowers. It is important to incorporate the individual–ESE score mechanism into the joint liability model as this can benefit both lenders and borrowers. For lenders, this mechanism provides a better understanding of the borrower’s risk profile, enabling them to make informed lending decisions and, in the process, ensure more sustainable joint liability contracts. It also provides information about the optimal group size for lending purposes. For borrowers, the mechanism provides them with a clearer picture of their ranking relative to their peers, allowing them to identify strengths and weaknesses and work on improving their individual–ESE scores. It also provides information about the number of peers needed for a group borrowing purpose, helping borrowers form groups that are more likely to succeed and have improved access to credit. Therefore we argue that incorporating the individual–ESE score mechanism into credit evaluation processes can lead to better–informed lending decisions, improved access to credit, and sustainable development outcomes.
@@
The remaining parts of this paper are organized as follows. We introduce the set-up and assumptions of the model in Section 2. We derive a set of results from the models consisting of a group size of two and a group size of $n$ in Sections 3 and 4 respectively. In Section 5 we refine the utility functional form of the model to that of a mean variance, which incorporates both the expected value of return and risk simultaneously in the face of climate change. As the resulting model does not have clean and closed-form analytical expressions, a set of simulation exercises is provided based on a set of climate scenarios. Section 6 concludes the paper with several qualifications and avenues for future research.
@@ 2 The Framework for an Individual ESE-Joint Liability Model
We are interested in examining a microcredit lending scenario in which a benevolent lender extends loans to a group of self–selected micro–entrepreneur borrowers who are jointly liable for repayment of the loan. Although each member of the group is given an individual loan, the group is collectively responsible for repaying the loans, and if the total loan is repaid, the entire group can qualify for a subsequent loan. The loans are then invested in a set of projects, which have an equal chance of success and can be either disjoint or correlated. It is assumed that the group members know the realized output of each project, while the lender does not. Both the lender and the borrowers have to make decisions in this set up. The lender has to decide on the optimal contract structure, the degree of joint liability, and the level of punishment in the event of a strategic default, while the borrowers have to decide whether to repay or default the loan.
@@
There are two prior studies which have inspired us in particular to write this paper. First rezaei2017optimal show that a group size can affect lending efficiency. Larger groups are more reliable in repaying loans for riskier projects, but the group size cannot become too large. Joint liability has a positive effect on borrowers’ repayment amount and welfare compared to individual liability, provided that group members impose strong punishment strategies against any strategically defaulting member. rezaei2017optimal also examine the possibility of relaxing some of the simplifying assumptions made in the basic model of joint liability. Second, in the context of climate change, castaing2021joint investigates the impact of risk–sharing networks on the adoption of weather shock management strategies in Burkina Faso; particularly, he studies how the principle of forced solidarity may reduce efforts to adopt techniques that mitigate exposure to climate change. The study finds that a system based on mutual assistance between farmers may reduce their incentives to adopt risk–mitigating strategies, both incrementally and transformationally. The study highlights the need for complementary mechanisms, such as insurance, to help farmers reduce their vulnerability to climate change.
@@
This paper pursues two primary objectives. The first one is to determine an optimal group size that maximizes the benefit for the borrowers, while ensuring that the lender breaks even. A larger group size can have a positive effect on the expected repayment amount, as more people are liable for repaying defaulted payments, thus ensuring a higher rate of repayment. However, a larger group size can also pose a threat to repaying members, as they must repay all defaulting peers’ repayments, when everyone else in the group defaults on repayment. Therefore this paper proposes a method to determine the optimal group size that enables joint liability to achieve its maximum loan ceiling. A second objective of this paper is to derive an optimal individual–ESE score. The importance of such a scoring system for both lenders and borrowers cannot be overstated. For lending companies, having access to a specific scoring mechanism for individual borrowers can provide a better understanding of their creditworthiness and repayment prospect. This information is essential for lending companies in determining the risk associated with lending to a particular borrower. Additionally the optimal group size of borrowers can be determined using the scoring system, which enables the lending company to mitigate risk by lending to a diversified portfolio. For borrowers, having a scoring system can equip them with a clearer picture of their ranking and eligibility for loans. This information is critical for farmers as they search for a group for the borrowing purposes since this can help them determine the number of peers needed in a group to secure a loan and improve their chances of obtaining a funding. Therefore, both lenders and borrowers can benefit from the implementation of such a comprehensive scoring system.
@@ 2.1 Model Assumptions
Prior to constructing the individual ESE–joint liability model, we state a number of assumptions underlying it.
@@
ASSUMPTION 1. The borrowers (who are the individual farmers) produce an agricultural product. For each borrower, the probability of high production (i.e., a successful project) is the same as given by $e$ and the probability of low production (i.e., a failed project) is given by $1-e$ respectively.
@@
ASSUMPTION 2. A higher individual–ESE score means that the farmer has a stronger ability to succeed in their project due to a high-quality comprehensive evaluation. For simplicity the probability of high yield is assumed to be linearly related to the borrower’s $ESE$ score; i.e., we assume that there is a linear relationship between $e$ and $ESE$ score: $e=k\cdot ESE+b$. The parameter $k$ represents the weight or strength of the impact of the borrower’s $ESE$ score on the probability of high yield. A higher value of $k$ indicates that a higher $ESE$ score leads to a more significant positive impact on the probability of success. This aligns well with the idea that a borrower’s positive $ESE$ practices contribute to a project success. In other words, higher $ESE$ scores indicate a better alignment with responsible practices. The constant $b$ represents a baseline/constant contribution of factors beyond an individual’s control, such as climatic and geographical conditions. Even if a borrower has a lower $ESE$ score, this constant term acknowledges that external factors can still influence the project’s success. It is essential to account for such factors in order to avoid unfairly penalizing borrowers for circumstances beyond their own control.
@@
The selection of the linear equation $e=k\cdot ESE+b$ is motivated by both the characteristics of the variables involved and its practical applicability. Notably, the non-negativity of $k$, $b$, and $ESE$ ensures that the resulting probability, denoted as $e$, maintains its intrinsic non-negative nature. As $ESE$ encompasses a continuous range from 0 to 100, it inherently aligns with this non-negative property. It is important to stress that the intuitive requirement for $e$ to lie between 0 and 1, which is a fundamental property of a probability, is elegantly met by setting $b=0$ and $k=\frac{1}{100}$. By adhering to these values, the probability of $e$ adheres to the established probability range. In addition, the model accommodates real-world scenarios by allowing for a non-zero baseline influence ($b>0$), often linked to exogenous factors. To this end, a modest value such as $b=0.3$ can be accommodated without violating the probability constraints. Notably, this adjustment merely necessitates setting $k=\frac{0.7}{100}$, thereby guaranteeing that $e$ remains within the bounds of a probability as it should be. While, at the same time, it also acknowledges that $e$ cannot span the interval from 0 to 0.3, this limitation is justified by underscoring $e$ as a representation of a probability of a high yield, which is inherently a positive entity in practical settings. Hence the rationale underlying the proposed equation effectively marries mathematical coherence with the nuanced demands from a real–world application.
@@
ASSUMPTION 3. The amount of loan secured by each borrower is denoted as $L$, while the amount for this borrower to be repaid is denoted as $w$. The unit of a selling price of the product is denoted as $p$, and the total income of a high production is denoted as $p\overline{Y}$, while the total income of a low production is denoted as $p\underline{Y}$.
@@
ASSUMPTION 4. High-yield borrowers repay their loans and low-yield borrowers pay back all the income they earn. As long as this group repays the total loans, this group can continue to secure subsequent loans. In the event that all borrowers are unsuccessful in their projects, the group will not be eligible for any future loans. This may result in a negative credit history for all members in the group, which could lead to further restrictions placed on them, such as being unable to use public transportation or purchase high-cost items. Therefore, if all borrowers fail, they may have to use all of their earnings to avoid receiving negative credit reports, even though their total earnings will be less than the total repayments.

### v3

@@ Abstract
This paper extends the application of ESG score assessment methodologies from large corporations to individual farmers’ production, within the context of climate change. Our proposal involves the integration of crucial agricultural sustainability variables into conventional personal credit evaluation frameworks, culminating in the formulation of a holistic sustainable credit rating referred to as the Environmental, Social, Economics (ESE) score. This ESE score is integrated into theoretical joint liability models, to gain valuable insights into optimal group sizes and individual-ESE score relationships. Additionally, we adopt a mean-variance utility function for farmers to effectively capture the risk associated with anticipated profits. Through a set of simulation exercises, the paper investigates the implications of incorporating ESE scores into credit evaluation systems, offering a nuanced comprehension of the repercussions under various climatic conditions.
@@ 1 Introduction
In recent years, the importance of incorporating sustainability factors into financial decision-making has gained significant traction across various sectors of the economy, particularly within agriculture and finance. While Environmental, Social, and Governance (ESG) metrics have become a cornerstone for evaluating corporate sustainability, a notable gap persists in assessing individual borrowers—particularly farmers seeking credit to finance their agricultural operations. To address this gap, this paper introduces a novel concept: a scoring system coined as Environmental, Sustainability, and Economics (ESE). The ESE score adapts the ESG framework to the agricultural sector, emphasizing economic considerations alongside environmental and social aspects as critical components of sustainable farming practices. This approach reflects the unique realities of individual farmers, where governance metrics are less applicable, and financial sustainability is paramount.
The ESE scoring system is designed to address the needs of multiple stakeholders who stand to benefit from such a framework. Policymakers and regulators can utilize ESE scores to design and implement targeted subsidies or tax incentives that promote sustainable farming practices. Financial institutions, such as banks and insurers, could leverage ESE metrics to assess the creditworthiness and risk profiles of individual farmers, while aligning their lending portfolios with sustainability goals. Similarly, agribusiness corporations can use ESE scores to evaluate their suppliers’ sustainability performance, ensuring greater transparency and alignment with their ESG-related commitments. Most importantly, farmers and cooperatives can benefit directly from the ESE scoring system, as it provides a structured framework to benchmark their practices, improve access to sustainable financing, and signal their commitment to sustainability to potential partners and stakeholders.
Although the ESE score draws inspiration from the ESG framework, there are both shared characteristics and critical differences between the two concepts. Both ESE and ESG scores emphasize sustainability and evaluate entities based on environmental and social metrics, such as resource use efficiency, emissions reduction, and biodiversity impacts. However, while ESG scores are primarily designed for corporations, investors, and financial markets, ESE scores are tailored specifically for the agricultural sector, focusing on individual farmers. A key distinction lies in the explicit inclusion of economic metrics in the ESE score, addressing the financial realities of farming—such as profitability and cost-efficiency—which are not standalone components in ESG. Additionally, ESG scores encompass broader social and governance considerations, such as diversity, labor practices, and organizational governance. In contrast, the ESE score narrows its scope to factors directly relevant to sustainable agriculture and individual creditworthiness. This granularity makes ESE scores a more specialized tool for supporting agricultural stakeholders while complementing broader ESG frameworks.
@@
In this paper, we also propose a novel link between individual ESE scores and a joint liability model, providing lenders and borrowers (where lenders are insurance companies and borrowers are farmers) with optimal scores and corresponding optimal group sizes that maximize the total utility of a group of individual borrowers. We begin our analysis with the simplest case of two borrowers in a group and later generalize it to a group of $n$ individual borrowers. Incorporating the individual ESE score mechanism into the joint liability model offers benefits to both lenders and borrowers. For lenders, this mechanism provides a better understanding of the borrower’s risk profile, enabling informed lending decisions and ensuring more sustainable joint liability contracts. Additionally, it offers insights into optimal group size for lending purposes, enhancing the efficiency of group-based borrowing. For borrowers, this mechanism offers a clearer picture of their ranking relative to peers, allowing them to identify areas for improvement and enhance their individual ESE scores. It also helps borrowers determine the ideal number of peers needed to form a group for borrowing purposes, improving group success rates and access to credit. Therefore, integrating the individual ESE score mechanism into credit evaluation processes can result in better-informed lending decisions, improved credit access, and enhanced sustainable development outcomes.
@@
The remaining parts of this paper are organized as follows. We list the steps and potential metrics for building an individual-ESE score mechanism in Section 2. In Section 3, we introduce the set-up and assumptions of the model and derive a set of results from the models consisting of a group size of two and a group size of $n$. In Section 4, we refine the utility functional form of the model to that of a mean-variance one, which incorporates both the expected value of return and risk simultaneously in the face of climate change. As the resulting model does not have cleanly closed-form analytical expressions, a set of simulation exercises is provided in this paper based on a set of climate scenarios. Section 5 concludes this paper.
2 The Mechanism of Individual-ESE Scores
In order to maintain accountability for sustainable agriculture commitments and provide valuable insights for policymaking, conducting regular and transparent assessments is of paramount importance. However, there are a number of variations in the definitions of sustainable agriculture, and also a limited number of quantitative assessments of agricultural sustainability that are currently available, according to Agovino et al. (2018). Although some experts view sustainable agriculture as a collection of management strategies, others see it as an ideology or a specific set of goals. Despite this, there is an increasing consensus that sustainable agriculture should be evaluated based on its environmental, economic, and social impacts, as emphasized by Zhang et al. (2021). Several frameworks and indicators have been developed. Existing sustainable agriculture indicators mainly assess the environmental impacts of agriculture production, and many have limited data availability. Although various organizations have called for monitoring agriculture globally, actual datasets to enable trend assessments are still currently lacking. The absence of a consistent quantification of agricultural sustainability across multiple dimensions impedes the successful identification of undesirable trade-offs and the development of win-win solutions. Given these challenges, relying on literature-based metrics is a reasonable starting point in developing a sustainable agricultural credit scoring system. Literature-based metrics are metrics that have been previously used and validated in other studies or publications and are therefore based on a larger sample size and a more diverse set of variables than any single data collection effort. Having literature-based metrics about agricultural credit scoring systems allows us to apply them when we obtain real data or use them to create a scoring system by using synthetic data. By compiling these synthetic metrics and adjusting them to fit the specific needs of the agricultural sector, we can create a credit scoring system that is more comprehensive and accurate than any single data collection effort.
The ultimate objective of this research is to eventually establish a comprehensive and sustainable agricultural credit scoring system at the individual level, which encompasses three critical aspects: Economic, Social, and Environmental. This system will be called an individual-ESE score. While traditional credit scoring systems focus mainly on borrowers’ economic and financial status, our system will take into account a broader range of factors that contribute to sustainability. We recognize the importance of economic metrics and acknowledge that they are crucial to the classical credit scoring system. Therefore, we will use some of the commonly used economic metrics in the related literature as a foundation for our economic metrics. Presented below are synopses of selected academic papers that have conducted research on the agricultural credit scoring system. (1) Römer and Mußhoff (2018) investigated the use of weather data in agricultural credit scoring for microfinance institutions, finding that simple correlation analysis between weather events and loan repayment is insufficient to forecast future repayment behavior. (2) Turvey (1991) reviewed the use of credit scoring for agricultural loans and its application to agricultural finance, identifying the importance of accurate data and the need for specialized credit scoring models for agricultural lending. (3) Lufburrow et al. (1984) developed a credit scoring model for farm loan pricing, utilizing statistical methods to determine the significant factors affecting loan repayment and default rates. (4) Limsombunchai et al. (2005) analyzed credit scoring for agricultural loans in Thailand, identifying the factors that affect credit scores and their impact on loan approval and interest rates. (5) Bumacov et al. (2014) examined the use of credit scoring in microfinance institutions, reporting that it can improve the outreach and efficiency of microfinance operations, but also highlighting the need for effective risk management strategies. (6) Mutamuliza et al. (2021) investigated the determinants of smallholder farmers’ participation in microfinance markets in Rwanda, identifying factors such as gender, education, and access to information that can affect farmers’ ability to access credit. We provide a summary of the key metrics employed in their credit scoring models as the foundations of economic metrics in Table 1.
Zhang et al. (2021) presents a variety of useful metrics for the environmental and social pillars. With additional information provided in Fuglie et al. (2016) and Movilla-Pateiro et al. (2021), we have compiled a comprehensive list of potential metrics for these pillars, along with detailed descriptions and methods of measurement. The metrics are shown in Table 2.
Metric Description Used as a Metric in the Scoring System
Age Age of the applicant in years (1), (4)
Gender Male, Female, Others (1), (6)
Education Education level of the applicant (6)
Collateral Collateral in thousands (1), (3), (4)
Assets Assets in thousands (1), (2), (3), (4), (5)
Debt Debt to other banks in thousands (1), (2), (3), (4), (5)
Deposit Deposit in the bank account in thousands (1), (2), (5)
Income Monthly business and household income in thousands (1), (2), (3), (4), (6)
Marital status Single, Married, Divorced, Others (1)
No. family members Number of family members (1)
Loan Size Loan amount in thousands (1), (4), (6)
Loan duration Loan Duration in Months (4)
Repayment ability Projected net cash flow plus projected inventory divided by a total line of credit (1), (2), (3)
Repayment history The average of loan principal repaid divided by principal due over the past three years (1), (3)
Tenure The ratio of acres owned and farmed to total acres farmed (3), (6)
Resident 1 if applicant is a resident; 0 otherwise (1),
Working experience Working experience in current profession in months (1),
Region and Country Region and Country of the Agricultural Business (1), (2), (4), (6)
Distance Distance to the microfinance institution in km (6)
Table 1: A List of Economic (E) Metrics Applied in Literature
Developing a comprehensive agricultural credit scoring system requires careful consideration of the various metrics that influence creditworthiness. Below is a scheme for generating a composite score based on multiple metrics, including how to assign weights to each metric. We aim at eventually developing a comprehensive ESE scoring system by either obtaining real data or generating synthetic data according to this scheme in the future.
  1. 1.
    Identify and gather relevant metrics: Begin by identifying the metrics that are most critical to assessing creditworthiness in agricultural lending, such as past repayment history, crop yields, farm management practices, and market conditions.
  2. 2.
    Determine the relative importance of each metric: Assign weights to each metric based on its perceived importance in predicting creditworthiness. This can be done through a combination of expert judgment and statistical analysis of historical data. Weights may also be assigned based on the level of risk exposure associated with each ESE factor, with higher weights assigned to metrics that pose a greater risk to a farmer’s performance. Notable, the overall Thomson Reuters ESG Score is calculated by applying an automated, factual logic that determines the weight of each category. Each category consists of a different number of measures. The count of measures per category determines the weight of the respective category.
  3. 3.
    Normalize the metrics to ensure that each has the same scale, range, and units of measurement. This can be achieved through methods such as standardization or normalization.
  4. 4.
    Once the weights and normalized values of each metric have been determined, calculate the composite score by multiplying the weight of each metric by its normalized value and summing the products.
  5. 5.
    Validate the composite score by comparing it to actual default rates or repayment history. Refine the weights and metrics as needed based on the validation results.
  6. 6.
    Continuously monitor and update the credit scoring system by incorporating new data and adjusting weights and metrics as needed to improve the accuracy and predictive power of the model.
  7. 7.
    By following the above steps, a comprehensive agricultural credit scoring system can be developed that considers multiple metrics and assigns weights based on their relative importance in predicting creditworthiness.
Sector Metric How to Measure
E Soil health Measuring the quality and health of the soil, including factors such as nutrient levels, organic matter content, and erosion rates. Soil testing kits could be used to measure.
E Water Consumption Modelled evapotranspiration using climate data (adjusted for crop type and water availability). Water usage tracking software could be used as a tool to measure.
E Water Quality N & P concentration (measured in stream, or modeled using land cover for landscapes)
E Habitat Conversion Remotely sensed land use (% of study area covered by natural habitat)
E Habitat Composition Remotely sensed land use (no. of land cover classes, indicating habitat diversity)
E Habitat Connectivity Remotely sensed land use (calculated connectivity score)
E Species Richness Field samples (richness relative to reference natural landscape)
E Species Abundance Field samples (abundance relative to reference natural landscape, i.e. the most undisturbed nearby similar habitat)
E Yield / Area Harvested mass of crop per unit area over a 5-year rolling average, perhaps using geometric mean to penalize variation in yield.
E Use of pesticides and chemicals Liters of pesticide or chemicals used per hectare
E Carbon Footprint The carbon footprint for different crop types are shown in Sah and Devakumar (2018). Measuring the amount of greenhouse gas emissions generated from the farmer’s activities, including fertilizer use, transportation, and energy use.
S Health & safety issues Whether the applicant ever had safety issues while participating in agricultural activities (0 if no, 1 if yes).
S Insuring health Whether the applicant has personal accident insurance or commercial insurance (0 if no, 1 if yes).
S Labor practices Measuring the fairness and safety of labor practices on the farm, including minimum wage compliance, safety regulations, and worker satisfaction.
S Community engagement Measuring the level of engagement and support the farmer provides to the local community, including economic benefits, charitable contributions, and community outreach.
S Ethical business practices Measuring the farmer’s adherence to ethical business practices, including transparency, honesty, and accountability.
S Compliance Measuring the farmer’s compliance with local and national regulations, including environmental, labor, and health and safety regulations.
Table 2: A list of Environmental (E) and Social (S) Metrics
@@ 3 The Framework for an Individual ESE-Joint Liability Model
We are interested in examining a microcredit lending scenario in which a benevolent lender extends loans to a group of self-selected micro-entrepreneur borrowers who are jointly liable for repayment of the loan. Although each member of the group is given an individual loan, the group is collectively responsible for repaying the loans, and if the total loan is repaid, the entire group can qualify for a subsequent loan. The loans are then invested in a number of projects, which have an equal chance of success and can be either disjoint from, or correlated with, each other. It is assumed that the group members know the realized output of each project, while the lender does not. Both the lender and the borrowers have to make decisions in this setup. The lender has to decide on the optimal contract structure, the degree of joint liability, and the level of punishment in the event of a strategic default, while the borrowers have to decide whether to repay or default, the loan.
@@
There are two prior studies which have inspired us in particular to write this paper. First Rezaei et al. (2017) show that a group size can affect lending efficiency. Larger groups are more reliable in repaying loans for riskier projects, but the group size cannot become too large. It is shown that joint liability has a positive effect on borrowers’ repayment amount and welfare compared to individual liability, provided that group members impose strong punishment strategies against any strategically defaulting member. In addition, Rezaei et al. (2017) also examine the possibility of relaxing some of the simplifying assumptions made in the basic model of joint liability. Second, in the context of climate change, Castaing (2021) investigates the impact of risk–sharing networks on the adoption of weather shock management strategies in Burkina Faso; particularly, she studies how the principle of forced solidarity may reduce efforts to adopt techniques that mitigate exposure to climate change. The study finds that a system based on mutual assistance between farmers may reduce their incentives to adopt risk–mitigating strategies, both incrementally and transformationally. The study highlights the need for complementary mechanisms, such as insurance, to help farmers reduce their vulnerability to climate change.
@@
Following the aforementioned two studies, this paper pursues two primary objectives. The first one is to determine an optimal group size that maximizes the benefit for the borrowers, while ensuring that the lender breaks even. A larger group size can have a positive effect on the expected repayment amount, as more individuals are liable for repaying defaulted payments, thus ensuring a higher rate of repayment. However, a larger group size can also pose a threat to repaying members, as they must repay all defaulting peers’ repayments, when everyone else in the group defaults on repayment. Therefore this paper proposes a method to determine the optimal group size that enables joint liability to achieve its maximum loan ceiling. A second objective of this paper is to derive an optimal individual–ESE score. The importance of such a scoring system for both lenders and borrowers cannot be overstated. For lending companies, having access to a specific scoring mechanism for individual borrowers can provide a better understanding of their creditworthiness and repayment prospects. This information is essential for lending companies in determining the risk associated with lending to a particular borrower. Additionally, the optimal group size of borrowers can be determined using the scoring system, which enables the lending company to mitigate risk by lending to a well-diversified portfolio. For borrowers, having a scoring system can equip them with a clearer picture of their ranking and eligibility for loans. This information is critical for farmers as they search for an ideal group for borrowing purposes since this can help them determine the number of peers needed in a group to secure a loan and, in the process, improve their chances of obtaining funding. Therefore, both lenders and borrowers can benefit from the implementation of such a comprehensive scoring system.
@@ 3.1 Model Assumptions
Below we state a number of assumptions underlying the proposed model.
@@
ASSUMPTION 1. The borrowers (who are the individual farmers) produce an agricultural product. For each borrower, the probability of high production (i.e., a successful project) is the same as given by $e$ and the probability of low production (i.e., a failed project) is given by 1e1-e respectively.
@@
ASSUMPTION 2. A higher individual–ESE score means that the farmer has a stronger ability to succeed in their project due to a high-quality comprehensive evaluation. For simplicity, the probability of high yield is assumed to be linearly related to the borrower’s ESE score; i.e., we assume that there is a linear relationship between $e$ and $ESE$: $e=k\cdot ESE+b$. The parameter $k$ represents the weight or strength of the impact of the borrower’s ESE score on the probability of high yield. A higher value of $k$ indicates that a higher ESE score leads to a more significant positive impact on the probability of success. This aligns well with the idea that a borrower’s positive ESE practices contribute to a project’s success. In other words, higher ESE scores indicate a better alignment with responsibly sustainable practices. The constant $b$ represents a baseline/constant contribution of factors beyond an individual’s control, such as climatic and geographical conditions. Even if a borrower has a lower ESE score, this constant term acknowledges that external factors can still influence the project’s success. Therefore it is essential to account for such factors to avoid unfairly penalizing borrowers for circumstances beyond their own control.
@@
The selection of the linear expression $e=k\cdot ESE+b$ is motivated by both the characteristics of the variables involved and its practical applicability. Notably, the non-negativity of $k$, $b$, and $ESE$ ensures that the resulting probability, denoted as $e$, maintains its intrinsic non-negative nature. As $ESE$ encompasses a continuous range from 0 to 100, it inherently aligns with this non-negative property. It is important to stress that the technical requirement for $e$, being a probability, to lie between 0 and 1, is conveniently met by setting $b=0$ and $k=\frac{1}{100}$. In addition, this specification also accommodates real-world situations by allowing for a non-zero baseline influence ($b>0$), often linked to exogenous factors. To this end, a modest value such as $b=0.3$ can be accommodated without violating the aforementioned probability constraint. Thus this adjustment necessitates setting $k=\frac{0.7}{100}$, thereby guaranteeing that $e$ remains within the bounds of a probability as required. While, at the same time, it also acknowledges that $e$ cannot span the interval from 0 to 0.3; this constraint is justified by underscoring that $e$ is a representation of a probability of a high yield, which is inherently a positive entity in practical settings. In summary, the rationale underlying the proposed equation effectively marries mathematical coherence with nuanced demands from real-world applications.
At first sight, the use of a sigmoid function to model $e$ in the model as $e=\frac{1}{1+\exp^{-ESE}}$ or $e=\frac{1}{1+\exp^{-(k\cdot ESE+b)}}$ appears to be an ideal choice. However, upon a closer examination, a critical observation related specifically to our study emerges; i.e., under the assumption that $k$, $b$, and $ESE$ are non-negative, the use of the sigmoid function would result in a constraint where $e$ precludes the values below $1/2$. This deviation from the alignment with realistic scenarios and the entire spectrum of probability values was deemed to be undesirable in the context of our study. Thus, while the use of the sigmoid function to model $e$ may possess certain merits in other contexts, its inherent restriction on precluding the probability values above $1/2$ in our model is at odds with the overarching goal of accurately representing the probability of high yield across its entire spectrum. As a result, a decision was made to retain the linear equation $e=k\cdot ESE+b$, which both adheres to the established criteria and provides the desired flexibility to represent a plausible range of probability values.
@@
ASSUMPTION 3. The amount of a loan secured by each borrower is denoted as $L$, while the amount for this borrower to be repaid is denoted as $w$. The unit selling price of the product is denoted as $p$, and the total income of a high production is denoted as $p\overline{Y}$, while the total income of a low production is denoted as $p\underline{Y}$.
@@
ASSUMPTION 4. High-yield borrowers repay their loans and low-yield borrowers pay back all of the income they earn. As long as this group repays the total loans, this group can continue to secure subsequent loans. In the event that all borrowers are unsuccessful in their projects, the group will not be eligible for any future loans. This may result in a negative credit history for all of the members in the group, which could lead to further restrictions placed on them, such as being unable to use public transportation or purchasing high-cost items. Therefore, if all borrowers fail in their projects, they may have to use all of their earnings to avoid receiving negative credit reports, even though their total earnings will be less than the total repayments.